Problem 2-9 Margin Calls on Short Sales (LO4, CFA5) You short sold 600 shares of stock at a price of $37 and an initial margin of 80 percent. If the maintenance margin is 40 percent, at what share price will you receive a margin call? What is your account equity at this stock price? (Do not round intermediate calculations. Round your answers to 2 decimal places.) Margin call price Account equity
Answer #1

The answer is:

Margin call price 47.57
Account equity 11,417.14

Calculations:

Proceeds from short sale = 600 shares*$37 = $22,200

Initial deposit = $22,200*80% = $17,760

Account value = 22200+17760 = $39,960

Thus margin call price = 39,960/[600+(0.40*600)] = 47.57

Account equity = 39,960 – (600*47.57) = 11,417.14